OpenStudy (anonymous):

Which is a rational number? a.square root of 8 b.square root of 16 c. square root of 18 d. square root of 20 Please give answer and explain. Thanks :D

OpenStudy (anonymous):

do you know what a rational number is

OpenStudy (anonymous):

No I do not

OpenStudy (anonymous):

number with no fraction, no?

OpenStudy (anonymous):

well if you find the square root of each number which is the only one that will square completely

OpenStudy (anonymous):

in other words \[\sqrt{a}\] is NOT a rational number unless \(a\) is a perfect square, like say 4,9,16,25,...

OpenStudy (anonymous):

ok so, only one squares completely so is that the answer?

OpenStudy (anonymous):

so for example \[\sqrt{25}\] is rational because \[\sqrt{25}=5\] but \[\sqrt{8}\] is not rational because 8 is not a perfect square

OpenStudy (anonymous):

Ok I get it thanks :D